The scores of individual students on a national achievement test have a normal distribution with a mean of 20.8 and a standard deviation of 1.9 at Feliciano delos Santos high school, 84 students took the test.if the scores of the students at this school have the same distribution as the national scores, how will you describe the sampling distribution of the scores of these students?assume that the population is infinite. Use the given steps above
1
Expert's answer
2021-05-07T14:10:01-0400

"\mu=20.8,\sigma=1.9,

n=84"


Using central limit theorem-


"\bar{X}" ~"~N(\mu,\dfrac{\sigma}{\sqrt{n}})"


So mean of sampling distribution "\bar{x}=\mu=20.8"


Standard deviation of samples "s=\dfrac{\sigma}{\sqrt{n}}=\dfrac{1.9}{\sqrt{84}}=0.2073"